OpenStudy (anonymous):

Calculate the volume of a sphere where the radius is 9 meters. 81 cubic meters 324 cubic meters 972 cubic meters 2916 cubic meters

OpenStudy (anonymous):

V≈3053.63m³

OpenStudy (anonymous):

You might wanna just go with 2916 cubic meters if that's the closest it gives you to 3053.63 cubic meters.

jimthompson5910 (jim_thompson5910):

The pi terms are missing from the answer choices (didn't paste properly)

jimthompson5910 (jim_thompson5910):

Use the formula \[\Large V = \frac{4}{3}\pi r^3\] in this case, r = 9